In Chromium(III) Chloride, CrCl3, chloride ions have cubic close packed arrangement and Cr(III) ions are present in the octahedral holes. What is the fraction of octahedral holes occupied? What is the fraction of total number of holes occupied?

Grade:upto college level

2 Answers

amit
7 years ago
in this questoin cr and cl dissociates in cr3+  and cl- when we draw the diagrham we see that in a cubic close structure one cromium ion play blance role with the 3 chloride ion therfore,,
cr3+=1 cl-=1/3
therefore 
in octahedral itwill be =n 
and
in tertrahedral void it will be=2*n
total voids=n+2n=3n
fraction =  n/3n=1/3
 
and therefore in ccp=1/3*1/3=1/9
therefore ratio will be 1/3:1/9
 
Sachin babar
6 years ago
In ccp there are no of atoms = 6 x 1/2 = 3No of atoms = no of octahedral voids = 3No of tetrahedral voids = 2 x3 = 6 and total nom of voids with chloride ions = 9 no of octahedral void occupied by Cr iii = 1fraction of octahedral void occupied = 1/3total fraction = 1/9